OpenStudy (anonymous):

Simplify. Please show all of your work. 3b – (-12 – b)

OpenStudy (anonymous):

Distribute the negative first then tshow me what you got

OpenStudy (anonymous):

3b - (12 -b)

OpenStudy (anonymous):

close there should be no parentheses and when distributing it looks something like this: -a(-b-c)= ab+ac

OpenStudy (anonymous):

Do that to your original equation

OpenStudy (anonymous):

3b*12+3b*b?

OpenStudy (anonymous):

not quite. Only distribute the negative one. For the moment forget about the 3b

OpenStudy (anonymous):

you should get 12+b do you see why?

OpenStudy (anonymous):

ok i think so

OpenStudy (anonymous):

Good now put the 3b back in so it looks like 3b+12+b you see?

OpenStudy (anonymous):

Now combine like terms so add the b's together

OpenStudy (anonymous):

3b+12?

OpenStudy (anonymous):

3b+b=?

OpenStudy (anonymous):

3b right

OpenStudy (anonymous):

no look at it like just adding the numbers in front of the b's (coefficients) so 3b+1b=? (b and 1b are the same thing)

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

oh so it is 4b

OpenStudy (anonymous):

4b+12

OpenStudy (anonymous):

up. :) plus 12. Good job. Thanks for hanging in there

OpenStudy (anonymous):

*yup

OpenStudy (anonymous):

no thank you

OpenStudy (anonymous):

No problem. If you could maybe just give me a medal im trying to get high leveled up. :p thanks